According to a Pew Research Center study, in May 2011, 32% of all American adults had a smart phone (one which the user can use to read email and surf the Internet). A communications professor at a university believes this percentage is higher among community college students. She selects 314 community college students at random and finds that 110 of them have a smart phone. Then in testing the hypotheses:

H0: p = 0.32 versus

Ha: p > 0.32,

what is the test statistic?

z =  . (Please round your answer to two decimal places.)
Answer #1

Solution:

Given

P0 = 32% = 0.32

1 - P = 1 - 0.32 = 0.78

n = 314

X = 110

= x / n = 110 / 314 = 0.350

z  =   - P0 /   (1 - P0 ) 198

z = 0.350 - 0.32 / ( 0.32 * 0.78 / 314 )

z = 1.15

This is the right tailed test